OpenStudy (anonymous):

|2x + 5| = 3x + 4 I got x = 1 and x = -9/5 and neither of those work. Is it possible to have no solutions?

OpenStudy (anonymous):

\(x=1\) should definitely work. Plugging in \(x=-\dfrac{9}{5}\), you have \[\left|2\left(-\dfrac{9}{5}\right)+5\right|=3\left(-\dfrac{9}{5}\right)+4\\ |\cdots|=-\frac{27}{5}+\frac{20}{5}\\ |\cdots|=-\frac{7}{5}\] Do you see why it doesn't work? Whenever you take the absolute value of something, you will never get a negative number.

OpenStudy (anonymous):

Oh gosh I wrote the wrong thing on my paper! xD Sorry thanks for pointing that out. I had a 4 instead of a 2 written down on accident. Thanks again!

OpenStudy (anonymous):

You're welcome!